Using transformation to graph a rational function
Explain how to use the graph of f(x)=1/x^2 to sketch the graph of g(x)=-1/(x-3)^2-9. Give the coordinates of the points on the graph of g that are found by transforming the points (-1,1) and (1,1) from the graph of f.
Please help me understand this question.

Answer by lwsshak3(11628)

Using transformation to graph a rational function
Explain how to use the graph of f(x)=1/x^2 to sketch the graph of g(x)=-1/(x-3)^2-9. Give the coordinates of the points on the graph of g that are found by transforming the points (-1,1) and (1,1) from the graph of f.
**
To sketch the graph of g from f, move the entire f curve 3 units to the right then drop it vertically 9 units. By this transformation, coordinates (-1,1) and (1,1) of f become (2,-8) and (4,-8) respectively in g.